OpenStudy (anonymous):

3. What is the 41st term of the arithmetic sequence where a1 = 15 and a15 = -41 ? help me! Im lost

OpenStudy (cp9454):

here first term is 15?

OpenStudy (anonymous):

yep, thats why I got messed up in the first place, it doesn't make sense

OpenStudy (cp9454):

a15= -41 a+14d = -41 15+14d = -41 14d= -56 d= -4 now, a= 15 and d= -4 a41 = a+40d = 15-40*4 = -145 Answer

OpenStudy (anonymous):

ok that makes sense now, but where did the "d" come from?

OpenStudy (cp9454):

15th term of series is given and 1st term is also given, so apply a15= a+14d put the value of a i.e. 1st term you will get eqn. in d only i.e. commen difference.

OpenStudy (anonymous):

Ohhhh, ok well thanks!